# Networked Multiagent Reinforcement Learning for Peer-to-Peer Energy Trading

### 25 Jan 2024 v1

@@ Introduction
The literature on decentralized control of energy transactions within distribution networks is extensive and diverse. Broadly speaking, there are two predominant approaches: the distributed optimization approach and P2P bilateral trading. One of the most prominent frameworks in the former approach is through the alternating direction of multiplier method (ADMM). The ADMM solves centralized economic dispatch/optimal power flow problems in a distributed manner by sending price signals (the dual variables of supply-demand constraints) for consumers/prosumers to independently optimize their solutions, which are then sent back to the central optimizer to update the price signals (for example, [6, 7]). While this iterative approach enjoys rigorous convergence guarantees under convexity, its practical application encounters notable challenges: (i) As an algorithmic approach rather than a market design, ADMM is subject to the same difficulties faced by a uniform-pricing-based market with all zero-marginal-cost resources. Specifically, it is prone to ‘bang-bang’ pricing issues – market clearing prices plummet to zero in oversupply or spike to a price ceiling in undersupply. Addressing this with additional market mechanisms, such as adding auxiliary and operating reserve markets, akin to wholesale markets, could overly complicate distribution-level markets for energy trading. (ii) ADMM and most of the distributed optimization algorithms are designed for static optimization, lacking the dynamism required for real-time adaptability, hence necessitating online versions of such algorithms. Online ADMM algorithms have been proposed in [8]. Even so, network constraints are only guaranteed at the algorithm convergence. How to guarantee the physical constraints in a distributed fashion within each iteration is an open question. (iii) The expectation that consumers or prosumers can solve optimization problems might be unrealistic due to a lack of expertise or resources.
@@
These limitations form the basis of our interest in exploring P2P bilateral trading. Consequently, our literature review will be primarily focused on this alternative approach. There have been multiple review papers on the large and ever-growing literature on P2P energy trading [9, 10, 11, 12, 13]. Within the literature, we focus on the works that explicitly consider distribution network constraints. [14] proposes a continuous double auction framework within which physical network constraints are maintained through a sensitivity-based approach. The agents in the auction are the so-called zero-intelligence-plus (ZIP) traders, who employ simple adaptive mechanisms without any learning or memory of the repeated multi-agent interactions. in which they propose sensitivity-based More sophisticated than ZIP, a fully decentralized MARL is adopted in [15] to realize decentralized voltage control, and it is extended in [2] to form a P2P energy trading market with physical constraints. These two papers implement a completely decentralized approach in the sense that each agent ignores the multi-agent interaction as well and just learns their single-agent policy through the Proximal Policy Optimization (PPO) algorithm [3] to maximize their own payoff.
@@
Another MARL approach that has been explored is the centralized training and decentralized execution (CTDE) framework. The multi-agent deep deterministic policy gradient (MADDPG) method from [4] was applied in [16] to solve the autonomous voltage control problem, and a multi-agent constrained soft actor-critic (MACSAC) algorithm inspired MADDPG is developed to perform Volt-VAR control (VVC). In [17], a MADDPG algorithm tailored for a double-sided auction market to realize P2P energy trading is proposed, though they do not consider physical constraints. While the CTDE framework potentially enables more efficient learning and robust outcomes, its reliance on a centralized entity for aggregating all agents’ information and coordinating their training introduces vulnerabilities. Specifically, it increases the risk of single-point failures and susceptibility to cyber-attacks.
@@
To overcome the drawbacks of fully decentralized and CTDE-based MARL, algorithms with limited communication within agents have also been developed. A notable contribution in this area is the consensus-based MARL algorithm designed for discrete-space problems, as presented in [18], and its extension to continuous-space problems in [1]. This model allows agents to learn and implement policies in a decentralized manner with limited communication, underpinned by a theoretical guarantee of convergence under certain conditions. Separately, [19] introduced a Scalable Actor Critic (SAC) framework, particularly relevant for large-scale networks where an individual agent’s state transition is influenced only by its neighbors. SAC is distinguished by its reduced communication needs compared to the CTDE framework and offers a theoretical guarantee for its policy convergence as well. However, the SAC framework is only established for discrete state space problems, and its extension and effectiveness in continuous-state space applications remain unexplored.
@@
In this work, we present several key contributions: First, we develop a comprehensive framework for a P2P energy trading market while integrating a consensus-based MARL algorithm. This framework facilitates automated control and bidding of DERs while allowing for decentralized learning of voltage constraints. Second, we present sufficient conditions required for the convergence of the MARL algorithm and show the details of implementing this algorithm using deep neural networks. Third, we conduct a comparative analysis of market outcomes derived from three distinct MARL strategies: a naive decentralization approach, the MADDPG approach, and the proposed consensus-based MARL. Our findings reveal that the consensus-based approach notably outperforms the others in achieving better market outcomes.
@@
In this section, we first describe a clearing mechanism, referred to as the supply-demand-ratio (SDR) mechanism [5], that can address the two potential issues faced by a renewable-dominated P2P market: (i) marginal-cost-based pricing yielding zero market prices most of the time, and (ii) the potential bang-bang outcomes in a double-auction-based mechanism as the reserve prices for buyers (the utility rates) and the sellers (usually the feed-in tariffs) are publicly known. The second issue is unique to the P2P energy market, in our opinion, and the bang-bang phenomenon is well-documented in [20] and will be explained further later in this section.
The SDR-clearing mechanism in P2P markets has been detailed in our earlier work [2]. For completeness, we briefly revisit it here. Let $\mathcal{I}={1,2,...,I}$ represent all consumers and prosumers, collectively termed as agents. We assume that trading among the agents takes place in fixed rounds (such as hourly or every 15 minutes) and is cleared ahead of time (such as day-ahead or hour-ahead). In each trading round $t$, each agent $i\in\mathcal{I}$ submits bids ($b_{i,t}$) to buy (if $b_{i,t}<0$) or sell (if $b_{i,t}>0$) energy. Agents can switch roles between buyers and sellers across different rounds, but not within the same round. For a round $t$, the sets of buyers and sellers are denoted as $\mathcal{B}_{t}=\{i:b_{i,t}<0\}$ and $\mathcal{S}_{t}=\{i:b_{i,t}\geq 0\}$, respectively.
@@
To ensure that $SDR_{t}$ is well-defined, we assume $\mathcal{B}_{t}\neq\emptyset$ for any $t$, a reasonable assumption given agents’ continuous need for energy. In rare cases where every agent is a prosumer with excess energy at time $t$, the P2P market can be temporarily suspended, and surplus energy sold to the grid at a predefined price, which does not impact our model or algorithm framework. Conversely, if $\mathcal{S}_{t}$ is $\emptyset$, indicating no excess energy for sale at $t$, then $SDR_{t}=0$.
Based on its definition, when $SDR_{t}>1$, it means that there is an oversupply. In this case, we assume that the excess energy in the P2P market is sold to a utility company, a distribution system operator (DSO), or an aggregator at a pre-defined rate, generally denoted as $FIT$ (feed-in tariff). [^1] The $FIT$ rate is a rate set by utilities/policymakers. It can be set as zero and will not affect the pricing mechanism in any way. If $10\leq SDR_{t}<$ 1, indicating overdemand in $t$, then the unmet demand bids will purchase energy at a pre-defined utility rate, denoted as $UR$. Without loss of generality, we assume that $FIT<UR$ [^2] If $FIT>UR$, it means that energy consumers pay more than the utility rate to purchase energy from the prosumers, which is equivalent to a direct subsidy from energy consumers (including low-income consumers who do not have the means to invest in DERs) to prosumers. This cannot be justified from an equity perspective.
@@
In (III), $P_{t}$ denote the market clearing price in round $t$, which is a piece-wise linear function with respect to the $SDR_{t}$, as illustrated in Figure 1.
@@
While the SDR approach may be criticized for not being based on economic theories or for being somewhat arbitrary (for example, instead of having a linear function, the pricing function could be replaced by any downward-shaped nonlinear functions, which would lead to different market prices with the same $SDR$), we want to point out that from our perspectives, the simplicity and transparency of the SDR mechanism outweigh the criticisms, and it is well-suited for a P2P energy market with all-zero marginal cost resources. In our previous work [20], it is identified that a P2P energy market using traditional double-sided auction may not work well due to the special characteristics of such a market. Since uncleared electricity demand bids have to be bought at the rate of $UR$, and uncleared energy supply bids have to be sold at $FIT$, with the $UR$ and $FIT$ known to all market participants. With all zero-marginal cost resources, agents would not know how to price their bids in a marginal-cost-based clearing mechanism, and the public information of the $UR$ and the $FIT$ naturally provides the focal points for sellers to bid at $UR$ and buyers to bid at $FIT$, aiming to maximize their own payoffs. This would lead to a similar bang-bang type market outcome: when there is overdemand, the market price would be UR; reversely, when there is oversupply, the price would be FIT. We view the SDR approach as a fair and reasonable alternative to the marginal-cost-based approach to avoid such bang-bang outcomes, especially for a new market in which the supply resources are likely much less than demand.
The SDR method offers two key benefits: First, it maintains the market clearing price between $UR$ and $FIT$, ensuring that transactions for buyers and sellers are at least as favorable as dealing directly with the utility. Second, it simplifies the bidding process by requiring only quantity bids. This allows consumers with flexible loads to concentrate on adjusting their demand timings and prosumers to better manage their energy storage and sales strategies, avoiding the complexity of submitting price-quantity pair bids.
@@
In this section, we first formulate the decision-making problem faced by each agent in a repeated P2P trading market as a Markov decision process. (For the sake of brevity, we only describe a prosumer’s problem. A pure consumer’s problem is the same without the supply bids.) We then introduce the details of the consensus-MARL algorithm.
@@
Due to the presence of energy storage, each agent’s bidding and charging decisions are linked over time, which are naturally modeled in a dynamic programming framework. Since an agent’s decisions at $t$ only depend on its energy storage level at $t-1$ and the PV generation in $t$, a discrete-time Markov decision process (MDP) model is suitable here. In the following, we describe the key building blocks of each agent’s MDP model.
@@
Observation and state variables: Agent $i\in\mathcal{I}$ at time step $t$ is assumed to receive the system states $s_{t}=(s_{1,t},\dots,s_{I,t})\in\mathcal{S}:=\Pi_{i=1}^{I}\mathcal{S}_{i}$, which is the concatenation of the personal states of each agent, where $I$ is the total number of agents in the system. Specifically, the personal states of agent $i$ are defined as $s_{i,t}:=(d^{p}_{i,t},d^{q}_{i,t},e_{i,t},PV_{i,t})\in\mathcal{S}_{i}$, where $d^{p}_{i,t}$ and $d^{q}_{i,t}$ are the inflexible demand of active and reactive power of agent $i$, $e_{i,t}$ is the state of charge of agent $i$’s energy storage system, and $PV_{i,t}$ is the PV (real) power generation in $t$.
@@
Among the state variables, $(d^{p}_{i,t},d^{q}_{i,t},PV_{i,t})$ are only observations because their transitions adhere to underlying distributions that are independent of agents’ actions. In contrast, the transitions of $e_{i,t}$ are influenced by agents’ actions, with details elaborated further below after agents’ actions are defined.
@@
The assumption that each agent is required to observe the whole system’s state may appear to be strong in a decentralized setting and is prone to criticisms and privacy concerns. Our responses are threefold: first, the identity of the agents is irrelevant, and hence, from an agent’s perspective, which state variable corresponds to which specific agent is not known. The state information can simply be labeled as Agent 1, 2, …, I, without revealing the agent’s identity. Second, the proposed algorithms are intended to be implemented on devices that can receive information from the grid (generally referring to a utility or a distributed system operator) and automatically participate in the repeated P2P market without human intervention (aka control automation). The information received by the devices should be encrypted and can only be decrypted by the control-automation devices. Granted that the devices could be compromised, there have been works on how to address adversarial attacks in a multi-agent reinforcement learning framework [21], which leads to interesting future research directions. Second, one of our main goals in this work is to establish rigorous convergence results for the proposed MARL algorithm. In [2], we proposed a completely decentralized framework in which each agent just uses their local state information to update the policies and solve their own MDP problem as if in a single-agent environment. The numerical results are surprisingly good in the sense that they all indicate convergence towards some steady states as the learning continues. However, no theoretical results could be established when each agent only uses local information, without communication or shared reward functions. Our view is that to be able to rigorously show a MARL game’s convergence to a steady state, there has to be some global information shared among the agents.
@@
Action or control variables: at time $t$, the actions that agent $i$ can take are represented by a vector $a_{i,t}:=(a^{q}_{i,t},a^{e}_{i,t})\in\mathcal{A}_{i}=\mathcal{A}^{q}_{i}\times\mathcal{A}^{e}_{i}$, where $a^{q}_{i,t}$ is the reactive power injected or absorbed by agent $i$’s smart inverter, and $a^{e}_{i,t}$ is energy charged (if $a^{e}_{i,t}>0$) to or discharged (if $a^{e}_{i,t}<0$) from the battery. The feasible action spaces of $a^{q}_{i,t}$ and $a^{e}_{i,t}$ are generically denoted by $\mathcal{A}^{q}_{i}$ and $\mathcal{A}^{b}_{i}$, respectively. In the specific context of our model, it is safe to assume that the action space $\mathcal{A}_{i}=\mathcal{A}^{q}_{i}\times\mathcal{A}^{e}_{i}$ is bounded. In addition, agent $i$’s actual bid (either buy or sell) to a P2P market at time $t$ can be written as
which is simply the net energy of PV generation minus baseload demand (of real power) and charge to the battery.
@@
where $\mathbb{I}_{i\in\mathcal{B}_{t}}$ is an indicator function such that $\mathbb{I}_{i\in\mathcal{B}_{t}}=1$ if $i\in\mathcal{B}_{t}$, and 0 otherwise. Similarly, $\mathbb{I}_{i\in\mathcal{S}_{t}}=1$ if $\in\mathcal{S}_{t}$ and 0 otherwise. $SDR_{t}$ is the supply-demand-ratio as defined in (1), which depends on all agents’ bids, and the bids further depend on each agent’s charging/discharging decisions, as well as the PV generation and the baseload, as specified in (3). $P_{t}$ in (6) is the market clearing price as defined in (III). Note that when $SDR_{t}<1$, not all demand bids will be cleared in the P2P market at time $t$. In such instances, they will need to purchase the needed energy at the utility rate, leading to the $UR\cdot b_{i,t}$ term when defining the reward $R_{i}^{m}$ in equation (6). To ensure fairness and prevent any inadvertent favoritism among the demand bids, we employ a mechanism where each demand bid is partially cleared. Specifically, the clearance proportion is exactly $SDR_{t}$ (which represents the percentage of total demand bids cleared), and the uncleared bid for every agent is then $(1-SDR_{t})\cdot b_{i,t}$. This approach of proportional scaling echoes the rules used in multi-unit double auctions in [22].
@@
The second term in the reward function, $R^{v}$, is the system voltage violation penalty at a given time $t$. To provide an explicit form of $R^{v}$, we first write out the standard bus-injection model for all $k=1,2,\dots,N$ (the time index $t$ is omitted for simplicity):
$\begin{split}p_{k}=&\sum_{j=1}^{N}|V_{k}||V_{j}|(G_{kj}\cos(\alpha_{k}-\alpha_{j})+B_{kj}\sin(\alpha_{k}-\alpha_{j})),\\ q_{k}=&\sum_{j=1}^{N}|V_{k}||V_{j}|(G_{kj}\sin(\alpha_{k}-\alpha_{j})-B_{kj}\cos(\alpha_{k}-\alpha_{j})).\end{split}$ (7)
@@
In the above system of nonlinear equations, $p_{k}$ and $q_{k}$, $k=1,\ldots,N$ are input data, where $p_{k}$ is the net real power injection (or withdrawal) at bus $k$ (summation of the energy amount bid from all agents at bus $k$ per unit time, i.e. $p_{k}=\sum_{i\in\mathcal{N}_{k}}b_{i}$, with $b_{i}$ given in (3) and $\mathcal{N}_{k}$ denoting the set of agents at bus $k$), and $q_{k}$ is the net reactive power flow at bus $k$ (i.e. $q_{k}=-\sum_{i\in\mathcal{N}_{k}}d^{q}_{i}+a^{q}_{i}$). The variables are $|V_{k}|$ and $\alpha_{k}$ for $k=1,\ldots,N$, with $|V_{k}|$ being the voltage magnitude and $\alpha_{k}$ being the phase angle at bus $k$. The terms $G_{kj}$ and $B_{kj}$ are parameters that represent the real and imaginary parts of the admittance of branch $k{\text{-}}j$ respectively.
@@
At a given round $t$ of the P2P trading, for a given set of the real and reactive power injection/withdrawal as the result of agents’ bidding $(p_{k},q_{k})_{k=1}^{N}$, if the system of equations (7) has a solution, [^3] Sufficient conditions under which (7) has a solution are provided in [23]. then we can define the voltage violation penalty as follows:
$R^{v}:=-\lambda\sum_{k=1}^{N}\text{clip}\Big{[}\max(|V_{k,t}|-\overline{V}_{k},\underline{V}_{k}-|V_{k,t}|),\{0,M\}\Big{]}.\vspace*{-5pt}$ (8)
@@
In (8), $\overline{V}_{k},\underline{V}_{k}$ represent the upper and lower limit of bus $k$’s voltage magnitude, respectively. Let $f_{clip}$ denote a generic ‘clip’ function in the form of $f_{clip}=\text{clip}[x,\{v_{min},v_{max}\}]$, which is defined as $f_{clip}=x$ when $v_{min}\leq x\leq v_{max}$, $f_{clip}=v_{min}$ if $x<v_{min}$, and $f_{clip}=v_{max}$ if $x>v_{max}$. Hence, if the voltage magnitudes $|V_{k,t}|$ obtained from solving equations (7) are within the voltage limit of $[\underline{V}_{k},\overline{V}_{k}]$ on every bus $k=1,\ldots,N$, then $R^{v}=0$; if there is voltage violation at a certain bus $k$, the term $\max(|V_{k,t}|-\overline{V}_{k},\underline{V}_{k}-|V_{k,t}|)$ becomes positive, and $R^{v}$ becomes a negative number with an arbitrary positive number $\lambda$ to amplify the voltage violation. The other positive parameter in (7), $M$, is just to ensure that the reward $R^{v}$ is bounded below, a technical condition needed to ensure the convergence of the consensus MARL algorithm to be introduced in the next section.
@@
If the system of equations (7) does not have a solution with a given set of agents’ bids, we can simply set $R^{v}=-\lambda\cdot N\cdot M$. We want to emphasize that the voltage violation does not result in actual financial penalties for the agents since the market clearing happens before real-time delivery. It simply represents feedback corresponding to agents’ bids and is used by each agent to train their reinforcement learning policies.

### 27 Jan 2024 v2

@@ Introduction
The literature on decentralized control of energy transactions in a distribution network is extensive and diverse. Broadly speaking, there are two predominant approaches: the distributed optimization approach and P2P bilateral trading. A key method in the former approach is the alternating direction multiplier method (ADMM), which solves centralized dispatch problems by iteratively updating price signals for consumer/prosumer optimization (such as [6, 7, 8]). Despite ADMM’s theoretical convergence under convexity, practical challenges arise: (i) As an algorithmic approach rather than a market design, ADMM is subject to the same difficulties faced by a uniform-pricing-based market of all zero-marginal-cost resources. Specifically, it is prone to ‘bang-bang’ pricing: market clearing prices plummet to zero in oversupply or spike to a price ceiling in undersupply. Addressing this with additional market mechanisms, such as adding ancillary service markets, akin to wholesale markets, could overly complicate distribution-level markets for energy trading. (ii) Most distributed algorithms, including ADMM, are static, lacking real-time adaptability. While online ADMM variants exist [9], ensuring network constraints within each iteration remains an open question. (iii) Assuming consumers or prosumers can solve complex optimization problems is often unrealistic due to potential expertise or resource limitations.
@@
These limitations form the basis of our interest in exploring P2P bilateral trading. Consequently, our literature review will focus on this alternative approach. There have been multiple review papers on the large and ever-growing literature on P2P energy trading [10, 11, 12, 13]. Within the literature, we focus on the works that explicitly consider distribution network constraints. [14] proposes a continuous double auction framework within which physical network constraints are maintained through a sensitivity-based approach. The agents in the auction are the so-called zero-intelligence-plus (ZIP) traders, who employ simple adaptive mechanisms without any learning of the repeated multi-agent interactions. More sophisticated than ZIP, a fully decentralized MARL is adopted in [15] to realize decentralized voltage control and is extended in [2] to form a P2P energy trading market with physical constraints. These two papers implement a completely decentralized approach in the sense that each agent ignores the multi-agent interaction as well and just learns their single-agent policy to maximize their own payoff.
@@
Another MARL approach that has been explored is the centralized training and decentralized execution (CTDE) framework, of which the multi-agent deep deterministic policy gradient (MADDPG) method [4] is a prominent example. It is applied in [16] to solve the autonomous voltage control problem In [17], a modified MADDPG algorithm is proposed for a double-sided auction market to facilitate P2P energy trading, although physical constraints are not considered in their model. While the CTDE framework can lead to more efficient learning and robust outcomes, its dependence on a centralized entity for collecting and coordinating all agents’ data introduces difficulties for practical implementations and potential vulnerabilities, such as increased risk of single-point failures and higher susceptibility to cyber-attacks
@@
To overcome the drawbacks of fully decentralized and CTDE-based MARL, algorithms with communication among agents have also been developed. A notable contribution in this area is the consensus-based MARL algorithm designed for discrete-space problems, as presented in [18], and its extension to continuous-space problems in [1]. This model allows agents to learn and implement policies in a decentralized manner with limited communication, underpinned by a theoretical guarantee of convergence under certain conditions. Separately, [19] introduced a Scalable Actor Critic (SAC) framework, particularly relevant for large-scale networks where an individual agent’s state transition is influenced only by its neighbors. SAC is distinguished by its reduced communication needs compared to the CTDE framework and offers a theoretical guarantee for its policy convergence as well. However, the SAC framework is only established for discrete state space problems, and its extension and effectiveness in continuous-state space applications remain unexplored.
@@
In this work, we present several key contributions: First, we develop a comprehensive framework for a P2P energy trading market while integrating a consensus-based MARL algorithm. This framework facilitates automated control and bidding of DERs while allowing for decentralized learning of voltage constraints. Second, we present the sufficient conditions required for the convergence of the MARL algorithm and show the details of implementing this algorithm using deep neural networks. Third, we conduct a comparative analysis of market outcomes derived from three distinct MARL strategies: a naive decentralization approach, the MADDPG approach, and the proposed consensus-based MARL. Our findings reveal that the consensus-based approach notably outperforms the others in achieving better market outcomes.
@@
In this section, we first describe a clearing mechanism, referred to as the supply-demand-ratio (SDR)[5], that can address the two potential issues faced by a renewable-dominated P2P market: (i) marginal-cost-based pricing yielding zero market prices, and (ii) the bang-bang outcomes in a double-auction-based mechanism, as the reserve prices for buyers (the utility rates) and the sellers (usually the feed-in tariffs) are publicly known. The second issue is unique to the P2P energy market, in our opinion, and the bang-bang phenomenon is documented in [20] and will be explained further in this section.
The SDR-clearing mechanism in P2P markets has been detailed in our earlier work [2]. For completeness, we briefly revisit it here. Let =$\mathcal{I}=\{1,2,...,I\}$ represent all consumers and prosumers, all referred to as agents. We assume that trading among the agents takes place in fixed rounds (such as hourly or every 15 minutes) and is cleared ahead of time (such as day-ahead or hour-ahead). In each trading round $t$, each agent $i\in\mathcal{I}$ submits bids ($b_{i,t}$) to buy (if $b_{i,t}<0$) or sell (if $b_{i,t}>0$) energy. Agents can switch roles between buyers and sellers across different rounds, but not within the same round. For a round $t$, the sets of buyers and sellers are denoted as $\mathcal{B}_{t}=\{i:b_{i,t}<0\}$ and $\mathcal{S}_{t}=\{i:b_{i,t}\geq 0\}$, respectively.
@@
To ensure that $SDR_{t}$ is well-defined, we assume $\mathcal{B}_{t}\neq\emptyset$ for any $t$, a reasonable assumption given agents’ continuous need for energy. In rare cases where every agent is a prosumer with excess energy at time $t$, the P2P market can be temporarily suspended, and surplus energy sold to the grid at a predefined price, which will not impact our model or algorithm framework. Conversely, if $\mathcal{S}_{t}$ is $\emptyset$, indicating no excess energy for sale at $t$, then $SDR_{t}=0$.
Based on its definition, when $SDR_{t}>1$, it means that there is an over-supply. In this case, we assume that the excess energy in the P2P market is sold to a utility company, a distribution system operator (DSO), or an aggregator at a pre-defined rate, generally denoted as $FIT$ (feed-in tariff). [^1] The $FIT$ rate is a rate set by utilities/policymakers. It can be set as zero and will not affect the pricing mechanism in any way. If $10\leq SDR_{t}<$ 1, indicating over-demand in $t$, then the unmet demand bids will purchase energy at a pre-defined utility rate, denoted as $UR$. Without loss of generality, we assume that $FIT<UR$ [^2] If $FIT>UR$, it means that energy consumers pay more than the utility rate to purchase energy from the prosumers, which is equivalent to a direct subsidy from energy consumers (including low-income consumers who do not have the means to invest in DERs) to prosumers. This cannot be justified from an equity perspective.
@@
In (III), $P_{t}$ denotes the market clearing price in round $t$, which is a piece-wise linear function with respect to the $SDR_{t}$, as illustrated in Figure 1.
@@
While the SDR approach may be criticized for not being based on economic theories, we argue that from our perspectives, the simplicity and transparency of the SDR mechanism outweigh the criticisms, and it is well-suited for a P2P energy market with all-zero marginal cost resources. In our previous work [20], it is identified that a P2P energy market using traditional double-sided auction may not work well due to the special characteristics of such a market. Since uncleared electricity demand bids have to be bought at the rate of $UR$, and uncleared energy supply bids have to be sold at $FIT$, with the $UR$ and $FIT$ known to all market participants. With all zero-marginal cost resources, agents would not know how to price their bids in a marginal-cost-based clearing mechanism, and the public information of the $UR$ and the $FIT$ naturally provides the focal points for sellers to bid at $UR$ and buyers to bid at $FIT$, aiming to maximize their own payoffs. This would lead to a bang-bang type market outcome: when there is overdemand, the market price would be UR; reversely, when there is oversupply, the price would be FIT. We view the SDR approach as a fair and reasonable alternative to the marginal-cost-based approach to avoid such bang-bang outcomes, especially for a new market in which the supply resources are likely much less than demand.
Additionally, the SDR method offers two key benefits: First, it maintains the market clearing price between $UR$ and $FIT$, ensuring that transactions for buyers and sellers are at least as favorable as dealing directly with the utility. Second, it simplifies the bidding process by requiring only quantity bids, avoiding the complexity of submitting bids in price-quantity pairs.
@@
In this section, we first formulate the decision-making problem faced by each agent as a MDP. (For brevity, we only describe a prosumer’s problem. A pure consumer’s problem is similar, just without the supply bids.) We then introduce the details of the consensus-MARL algorithm.
@@
Due to the presence of energy storage, each agent’s bidding and charging decisions are linked over time, which are naturally modeled through dynamic programming. Since an agent’s decisions at $t$ only depend on its energy storage level at $t-1$ and the PV generation in $t$, a discrete-time MDP model is suitable here. In the following, we describe the key building blocks of each agent’s MDP model.
@@
Observation and state variables: Agent $i\in\mathcal{I}$ at time step $t$ is assumed to receive the system states $s_{t}=(s_{1,t},\dots,s_{I,t})\in\mathcal{S}:=\Pi_{i=1}^{I}\mathcal{S}_{i}$, which is the concatenation of the states of each agent, where $I$ is the total number of agents. Specifically, the individual state variable is defined as $s_{i,t}:=(d^{p}_{i,t},d^{q}_{i,t},e_{i,t},PV_{i,t})\in\mathcal{S}_{i}$, where $d^{p}_{i,t}$ and $d^{q}_{i,t}$ are the inflexible demand of active and reactive power of agent $i$, $e_{i,t}$ is the state of charge of agent $i$’s energy storage system, and $PV_{i,t}$ is the PV (real) power generation in $t$.
@@
Among the state variables, $(d^{p}_{i,t},d^{q}_{i,t},PV_{i,t})$ are only observations because their transitions adhere to underlying distributions that are independent of agents’ actions. In contrast, the transition of $e_{i,t}$ is influenced by agents’ actions, with details to be elaborated further below after agents’ actions are defined.
@@
The assumption that each agent is required to observe the whole system’s state may appear to be strong in a decentralized setting and is prone to criticisms and privacy concerns. Our responses are twofold: first, the identity of the agents is irrelevant, and hence, from an agent’s perspective, which state variable corresponds to which specific agent is not known. Second, the proposed algorithms are intended to be implemented on devices that can receive information from the grid (generally referring to a utility or a distributed system operator) and automatically participate in the repeated P2P market without human intervention (aka control automation). The information received by the devices should be encrypted and can only be decrypted by the control-automation devices. Granted that the devices could be compromised, there have been works on how to address adversarial attacks in a multi-agent reinforcement learning framework [21], which leads to interesting future research directions.
@@
Action or control variables: at time $t$, the actions that agent $i$ can take are represented by a vector $a_{i,t}:=(a^{q}_{i,t},a^{e}_{i,t})\in\mathcal{A}_{i}=\mathcal{A}^{q}_{i}\times\mathcal{A}^{e}_{i}$, where $a^{q}_{i,t}$ is the reactive power injected or absorbed by agent $i$’s smart inverter, and $a^{e}_{i,t}$ is energy charged (if $a^{e}_{i,t}>0$) to or discharged (if $a^{e}_{i,t}<0$) from the battery. The feasible action spaces of $a^{q}_{i,t}$ and $a^{e}_{i,t}$ are generically denoted by $\mathcal{A}^{q}_{i}$ and $\mathcal{A}^{b}_{i}$, respectively. In the specific context of our model, it is safe to assume that the action space $\mathcal{A}_{i}=\mathcal{A}^{q}_{i}\times\mathcal{A}^{e}_{i}$ is bounded.
@@
where $\mathbb{I}_{i\in\mathcal{B}_{t}}$ is an indicator function such that $\mathbb{I}_{i\in\mathcal{B}_{t}}=1$ if $i\in\mathcal{B}_{t}$, and 0 otherwise. Similarly, $\mathbb{I}_{i\in\mathcal{S}_{t}}=1$ if $\in\mathcal{S}_{t}$ and 0 otherwise. $SDR_{t}$ is the supply-demand-ratio as defined in (1), which depends on all agents’ bids, and the bids further depend on each agent’s charging/discharging decisions, as well as the PV generation and the baseload. Specifically, the bids can be explicitly written as
which is simply the net energy of PV generation minus baseload demand (of real power) and charge to the battery. The market clearing price $P_{t}$ in (5) is defined in (III).
@@
Note that when $SDR_{t}<1$, not all demand bids will be cleared in the P2P market at time $t$. In such instances, they will need to purchase the needed energy at the utility rate, leading to the $UR\cdot b_{i,t}$ term when defining the reward $R_{i}^{m}$ in equation (5). To ensure fairness and prevent any inadvertent favoritism among the demand bids, we employ a mechanism where each demand bid is partially cleared. Specifically, the clearance proportion is exactly $SDR_{t}$ (which represents the percentage of total demand bids cleared), and the uncleared bid for every agent is then $(1-SDR_{t})\cdot b_{i,t}$. This approach of proportional scaling echoes the rules used in multi-unit double auctions in [22].
@@
The second term in the reward function, $R^{v}$, is the system voltage violation penalty at a given time $t$. To provide an explicit form of $R^{v}$, we first write out the standard bus-injection model. Assume that the network has $N$ buses, we have that for $\kappa=1,\ldots,N$ (the time index $t$ is omitted below for simplicity):
$\begin{split}p_{\kappa}=&\sum_{j=1}^{N}|V_{\kappa}||V_{j}|(G_{\kappa j}\cos(\alpha_{\kappa}-\alpha_{j})+B_{\kappa j}\sin(\alpha_{\kappa}-\alpha_{j})),\\ q_{\kappa}=&\sum_{j=1}^{N}|V_{\kappa}||V_{j}|(G_{\kappa j}\sin(\alpha_{\kappa}-\alpha_{j})-B_{\kappa j}\cos(\alpha_{\kappa}-\alpha_{j})).\end{split}$ (7)
@@
In the above system of nonlinear equations, $p_{\kappa}$’s and $q_{\kappa}$’s are input data, with $p_{\kappa}$ being the net real power injection (or withdrawal) at bus $\kappa$ (i.e. $p_{\kappa}=\sum_{i\in\mathcal{N}_{\kappa}}b_{i}$, with $b_{i}$ given in (6) and $\mathcal{N}_{\kappa}$ denoting the set of agents at bus $\kappa$), and $q_{k}$ being the net reactive power flow at bus $k$ (i.e. $q_{\kappa}=\sum_{i\in\mathcal{N}_{\kappa}}(-d^{q}_{i}+a^{q}_{i})$). The variables are $|V_{\kappa}|$ and $\alpha_{\kappa}$ for $\kappa=1,\ldots,N$, with $|V_{\kappa}|$ being the voltage magnitude and $\alpha_{\kappa}$ the phase angle at bus $\kappa$. The terms $G_{\kappa j}$ and $B_{\kappa j}$ are parameters that represent the real and imaginary parts of the admittance of branch $\kappa{\text{-}}j$ in the network, respectively.
@@
At a given round $t$ of the P2P trading, for a given set of the real and reactive power injection/withdrawal as the result of agents’ bidding, if the system of equations (7) has a solution, [^3] Sufficient conditions under which (7) has a solution are provided in [23]. then we can define the voltage violation penalty as follows:
$R^{v}:=-\lambda\sum_{\kappa=1}^{N}\text{clip}\Big{[}\max(|V_{\kappa,t}|-\overline{V}_{\kappa},\underline{V}_{\kappa}-|V_{\kappa,t}|),\{0,M\}\Big{]}.\vspace*{-5pt}$ (8)
@@
In (8), $\overline{V}_{\kappa}$ and $\underline{V}_{\kappa}$ represent the upper and lower limit of bus $k$’s voltage magnitude, respectively. Let $f_{clip}$ denote a generic ‘clip’ function in the form of $f_{clip}=\text{clip}[x,\{v_{min},v_{max}\}]$, which is defined as $f_{clip}=x$ when $v_{min}\leq x\leq v_{max}$, $f_{clip}=v_{min}$ if $x<v_{min}$, and $f_{clip}=v_{max}$ if $x>v_{max}$. Based on the clip function definition, if the voltage magnitudes $|V_{\kappa,t}|$ obtained from solving equations (7) are within the voltage limit of $[\underline{V}_{\kappa},\overline{V}_{\kappa}]$ on every bus $\kappa=1,\ldots,N$, then $R^{v}=0$; if there is voltage violation at a certain bus $\kappa$, the term $\max(|V_{\kappa,t}|-\overline{V}_{\kappa},\ \underline{V}_{\kappa}-|V_{\kappa,t}|)$ becomes positive, and $R^{v}$ becomes a negative number with an arbitrary positive number $\lambda$ to amplify the voltage violation. The other positive parameter in (7), $M$, is just to ensure that the reward $R^{v}$ is bounded below, a technical condition needed to ensure the convergence of the consensus MARL algorithm to be introduced in the next section.
@@
If (7) does not have a solution with a given set of agents’ bids, we can simply set $R^{v}=-\lambda\cdot N\cdot M$. We want to emphasize that the voltage violation does not result in actual financial penalties for the agents since the market clearing happens before real-time delivery. It is simply feedback sent to agents for them to train their reinforcement learning policies.